OpenStudy (anonymous):

which ordered pair is a solution of y = -6x-7? a.(0.5,-4) b.(2,-19) c.(-2,-5) d.(3,11)

OpenStudy (anonymous):

b

OpenStudy (anonymous):

how can you explain me how you get that?

OpenStudy (anonymous):

this can be solved by substitution of (x,y) for: a implies -4 = -6(0.5)-7 = -3-7 =-10, which is a contradiction. b implies -19 = -6(2)-7 = -12-7 =-19 True c implies -5 = -6(-2)-7 = 12-7 = 5, contradiction d implies 11 = -6(3)-7 =-18-7 =-25, contradiction

OpenStudy (anonymous):

b